OpenStudy (anonymous):

let f(x)=1-2x^2 and g(x) = x+1 compute the following (f o g)(x)

OpenStudy (anonymous):

\[f\circ g(x)=f(g(x))=f(x+1)=1-2(x+1)^2\] then algebra

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

now i have (f o G)(-1)

OpenStudy (anonymous):

do I just put the -1 in place of all the x's?